OpenStudy (anonymous):

Question 3 of 5 If f(x) varies directly with x2, and f(x) = 96 when x = 4, find the value of f(2). (Points : 1) a. 192 b. 72 c.32 d.24 which one, hurry. I just need the answer, I know how to do it, its just im stuck on this one. Please and thank you

OpenStudy (anonymous):

Someone help please

OpenStudy (anonymous):

is that x^2 or x2

OpenStudy (anonymous):

x2

OpenStudy (anonymous):

wait hold on let me check again

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

well we have f(x)= c * x^2 so f(4)= c* 4^2 = 96 then we got c=6 f(2)=6x2^2=24

OpenStudy (anonymous):

so its D

OpenStudy (anonymous):

thank you so much (:

OpenStudy (anonymous):

so welcome